OpenStudy (anonymous):

what is an equal equation to (2^3+ 3^2). plz help

OpenStudy (anonymous):

(2^3+ 3^2) 2^3 = 2*2*2 3^2 = 3*3 so (2^3+3^2) = ((2*2*2)+(3*3))

OpenStudy (amistre64):

question is wierd, what is an equal equation to 24?

OpenStudy (anonymous):

ok thanks tomo ..but they have answers like 6^6 & 5^5 & 4^2 + 1. ????

OpenStudy (amistre64):

soo, essentially, they are asking you to find the value of the given, and equate it to the value of an option?

OpenStudy (amistre64):

what is the value of 2^3 + 3^2 ?

OpenStudy (anonymous):

8+9=17 but its not an option.

OpenStudy (amistre64):

but, 17 IS an option which option also has a value of 17?

OpenStudy (anonymous):

6^6 5^5 4^2 + 1. ????

OpenStudy (anonymous):

Rite that's why im so confused... what's the correct answer ?

OpenStudy (anonymous):

4^2+ 1= 17????? @amistre 64

OpenStudy (anonymous):

correcttttt

OpenStudy (anonymous):

Thank you........Sooooo much guys!